# An iterative method for the solution of Laplace-like equations in high and very high space dimensions

### v3

@@ 1 Introduction
The numerical solution of partial differential equations in high space dimensions is a difficult and challenging task. Methods such as finite elements, which work perfectly in two or three dimensions, are not suitable for solving such problems because the effort grows exponentially with the dimension. Random walk based techniques only provide solution values at selected points. Sparse grid methods are best suited for problems in still moderate dimensions. Tensor-based methods Bachmayr , Hackbusch_1 , Khoromskij stand out in this area. They are not subject to such limitations and perform surprisingly well in a large number of cases. Tensor-based methods exploit less the regularity of the solutions rather than their structure. Consider the equation
@@
$U_{k+1}=\,U_{k}\,-\,(-\Delta+\mu)^{-1}(\mathcal{L}U_{k}+\mu U_{k}-F).$ (8)
The calculation of these iterates requires the solution of equations of the form (1), that is, the calculation of integrals of the form (2), now not over $\mathbb{R}^{m}$, but over the higher dimensional $\mathbb{R}^{n}$. The symbol $\|T^{t}\omega\|^{2}$ of the operator $\mathcal{L}$ is a homogeneous second-order polynomial in $\omega$. For separable right-hand sides $F$ as above, the calculation of these integrals thus reduces to the calculation of products of lower, in the extreme case one-dimensional integrals.
The reason for the usually astonishingly good approximation properties of these iterates is the directional behavior of the term $\|T^{t}\omega\|^{2}$, more precisely the fact that the euclidean norm of the vectors $T^{t}\eta\in\mathbb{R}^{m}$ for vectors $\eta$ on the unit sphere $S^{n-1}$ of the $\mathbb{R}^{n}$ takes an almost constant value outside a very small set whose size decreases rapidly with increasing dimensions, a typical concentration of measure effect. To capture this phenomenon quantitatively, we introduce the probability measure
@@
on the Borel subsets $M$ of the $\mathbb{R}^{n}$, where $\nu_{n}$ is the volume of the unit ball and $n\nu_{n}$ thus is the area of the unit sphere. If $M$ is a subset of the unit sphere, $\mathbb{P}(M)$ is equal to the ratio of the area of M to the area of the unit sphere. If $M$ is a sector, that is, if $M$ contains with a vector $\omega$ also its scalar multiples, $\mathbb{P}(M)$ measures the opening angle of $M$. In the following we assume that the mean value of the expression $\|T^{t}\eta\|^{2}$ over the unit sphere, or in other words its expected value with respect to the given probability measure, takes the value one. This is only a matter of the scaling of the variables in the higher dimensional space and does not represent a restriction. The decisive quantity is the variance of the expression $\|T^{t}\eta\|^{2}$, considered as a random variable on the unit sphere. Since the values $\|T^{t}\eta\|^{2}$ are, except for rather extreme cases, approximately normally distributed in higher dimensions, the knowledge of the variance makes it possible to capture the distribution of these values sufficiently well. The concentration of measure effect is reflected in the fact that the variances almost always tend to zero as the dimensions increase.
@@
with respect to a broad range of Fourier-based norms. Provided the values $\|T^{t}\eta\|^{2}$ are approximately normally distributed with expected value $E=1$ and small variance $V$, the measure (9) of the sector $S$ is almost one as soon as $\rho$ exceeds the standard deviation $\sigma=\sqrt{V}$ by more than a moderate factor. The sector $S$ then fills almost the entire frequency space. This is admittedly an idealized situation and the actual convergence behavior is more complicated, but this example characterizes pretty well what one can expect. The higher the dimensions and the smaller the variance, the faster the iterates approach the solution.
@@
In a sense, the following two sections form the core of the present work. They deal with the distribution of the values $\|T^{t}\eta\|^{2}$ on the unit sphere. Section 3 treats the general case. We first show that the expected value $E$ and the variance $V$ of the expression $\|T^{t}\eta\|^{2}$, considered as a random variable, can be expressed in terms of the singular values of the matrix $T$. Using this representation, we show that for random matrices $T$ with assigned expected value $E=1$ the expected value of the corresponding variances $V$ not only tends to zero as the dimension $m$ goes to infinity, but also that these variances cluster the more around their expected value the larger $m$ is. Sampling large numbers of these variances supports these theoretical findings. We also study a class of matrices $T$ that are associated with interaction graphs. These matrices correspond to the case that some or all coordinate differences are introduced as additional variables and formed the original motivation for this work. The expected values $E$ that are assigned to these matrices always take the value one and the corresponding variances $V$ can be expressed directly in terms of the vertex degrees and the dimensions. Numerical experiments for standard classes of random graphs show that the variances tend to zero with high probability as the number of the vertices increases. As mentioned, the knowledge of the expected value and the variance assigned to a given matrix $T$ is of central importance since the distribution of the values $\|T^{t}\eta\|^{2}$ is usually not much different from the corresponding normal distribution. This can be checked experimentally by sampling these values for a large number of points $\eta$ that are uniformly distributed on the unit sphere.
These observations are supported by the results for orthogonal projections from a higher onto a lower dimensional space presented in Sect. 4, a case which is accessible to a complete analysis and that for this reason serves us as a model problem. These results are based on a precise knowledge of the corresponding probability distributions and are related to the random projection theorem and the Johnson-Lindenstrauss lemma Vershynin . This section is based in some parts on results from our former papers Yserentant_2020 and Yserentant_2022 , which are taken up here and are adapted to the present situation.
In Sect. 5 we finally return to the higher-dimensional counterpart (7) of the original equation (1) and study the convergence behavior of the polynomially accelerated version of the iteration (8) for its solution. Special attention is given to the limit case $\mu=0$ of the Laplace equation. The section ends with a brief review of an approximation of the form (4) by sums of Gauss functions.
@@ 2 Solutions as traces of higher-dimensional functions
$\|U\|=\left(\frac{1}{\sqrt{2\pi}}\right)^{n}\!\int|\widehat{U}(\omega)|\,\mathrm{d}\omega.$ (2)
@@
$(\mathrm{D}^{\alpha}u)(x)=\left(\frac{1}{\sqrt{2\pi}}\right)^{n}\!\int(\mathrm{i}\,T^{t}\omega)^{\alpha}\widehat{U}(\omega)\,\mathrm{e}^{\,\mathrm{i}\,\omega\cdot Tx}\,\mathrm{d}\omega,$ (5)
which is, like $u$, itself uniformly continuous and vanishes at infinity.
@@
$|\,U|_{T}=\left(\frac{1}{\sqrt{2\pi}}\right)^{n}\!\int\|T^{t}\omega\|^{2}\,|\widehat{U}(\omega)|\,\mathrm{d}\omega.$ (6)
@@
From now on, the equation (9) will replace the original equation (11). Our aim is to compute its solution (10) iteratively by polynomial accelerated versions of the basic iteration (8). The convergence properties of this iteration depend decisively on the directional behavior of the values $\|T^{t}\omega\|^{2}$, which will be studied in the next two sections, before we return to the equation and its iterative solution.
@@
The norms split into two groups, both of which depend on a smoothness parameter. The norms in the first group possess the radial-angular representation
$\|U\|_{s}=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\bigg{(}n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r\bigg{)}\,\mathrm{d}\eta$ (15)
and take in cartesian coordinates the more familiar form
$\|U\|_{s}=\int_{\mathbb{R}^{n}}|\widehat{U}(\omega)|\|\omega\|^{s}\,\mathrm{d}\omega.$ (16)
The spectral Barron spaces $\mathcal{B}_{s}=\mathcal{B}_{s}(\mathbb{R}^{n})$, $s\geq 0$, consist of the functions $U$ with finite norms $\|U\|_{t}$ for $0\leq t\leq s$. Since all these norms scale differently, we do not combine them to an entity but treat them separately. One should be aware that these norms are quite strong. For integer values $s$, functions $U\in\mathcal{B}_{s}$ possess continuous partial derivatives up to order $s$, which are bounded by the norms $\|U\|_{k}$, $k\leq s$, and vanish at infinity. Barron spaces play a prominent role in the analysis of neural networks and in high-dimensional approximation theory in general; see for example Siegel-Xu .
The norms in the second group are a mixture between an $L_{1}$-based norm, in radial direction, and an $L_{2}$-based norm for the angular part and are given by
$|\!|\!|U|\!|\!|_{s}^{2}=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\bigg{(}n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r\bigg{)}^{2}\,\mathrm{d}\eta.$ (17)
If the norm $|\!|\!|U|\!|\!|_{s}$ of $U$ is finite, then also the norm $\|U\|_{s}$ and $\|U\|_{s}\leq|\!|\!|U|\!|\!|_{s}$ holds. This follows from the Cauchy-Schwarz inequality. Both norms scale in the same way. By (14), they coincide for radially symmetric functions.
The following lemma is needed to bring these two norms into connection with Theorem 2.1 and the solution (10) of the equation (9).
@@
Let $m\geq 3$ and $m\geq 5$, respectively. Then the surface integrals
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\frac{1}{\|T^{t}\eta\|^{2}}\,\mathrm{d}\eta,\quad\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\frac{1}{\|T^{t}\eta\|^{4}}\,\mathrm{d}\eta$ (18)
take finite values, despite the singular integrands.
@@
Let $\Sigma_{0}$ be the $(m\times m)$-diagonal matrix with the singular values of $T$ on its diagonal. A variable transform and the Fubini-Tonelli theorem then lead to
$\int_{\mathbb{R}^{n}}\frac{1}{\|T^{t}\omega\|^{2}}\,\mathrm{e}^{-\|\omega\|^{2}}\!\,\mathrm{d}\omega\;=\int_{\mathbb{R}^{m}}\frac{1}{\|\Sigma_{0}\omega_{1}\|^{2}}\,\mathrm{e}^{-\|\omega_{1}\|^{2}}\!\,\mathrm{d}\omega_{1}\int_{\mathbb{R}^{n-m}}\!\mathrm{e}^{-\|\omega_{2}\|^{2}}\!\,\mathrm{d}\omega_{2}.$
Because the first of the two integrals on the right-hand side of this identity is finite for dimensions $m\geq 3$ and the second without such a constraint, the integral
$\int_{\mathbb{R}^{n}}\frac{1}{\|T^{t}\omega\|^{2}}\,\mathrm{e}^{-\|\omega\|^{2}}\!\,\mathrm{d}\omega\;=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\frac{1}{\|T^{t}\eta\|^{2}}\,\mathrm{d}\eta\;\int_{0}^{\infty}\frac{1}{r^{2}}\,\mathrm{e}^{-r^{2}}r^{n-1}\,\mathrm{d}r$
takes a finite value. The same then also holds for the integral over the unit sphere on the right-hand side of the equation. For dimensions $m\geq 5$, the second of the two integrals in (18) can be treated in the same way. ∎
For a given function $F$ with integrable Fourier transform, the function
$\widehat{U}(\omega)=\frac{1}{\|T^{t}\omega\|^{2}}\,\widehat{F}(\omega)$ (19)
possesses the radial-angular representation
$\widehat{U}(r\eta)=\frac{1}{\|T^{t}\eta\|^{2}}\,\frac{\widehat{F}(r\eta)}{r^{2}}.$ (20)
If the function $\phi:S^{n-1}\to\mathbb{R}$ given by
$\phi(\eta)=n\nu_{n}\int_{0}^{\infty}\bigg{|}\frac{\widehat{F}(r\eta)}{r^{2}}\bigg{|}\,r^{n-1}\,\mathrm{d}r$ (21)
is bounded and the dimension $m$ is three or higher, the function (20) is integrable by Lemma 2 and Fourier transform of a function $U\in\mathcal{B}_{0}(\mathbb{R}^{n})$ with finite norm $\|U\|_{0}$. For dimensions $m$ greater than four, it suffices that $\phi$ is square integrable. For the other norms introduced above, one can argue in a similar way.
3 A moment-based analysis
The purpose of this and the next section is to study how much the pseudo-differential operator (7) differs from the negative Laplace operator, that is, how much
$\chi(\omega)=\|T^{t}\omega\|^{2}$ (1)
deviates on the unit sphere $S^{n-1}$ of the $\mathbb{R}^{n}$ from the constant one. The expression (1) is a homogeneous function of second order in the variable $\omega$. To study its angular distribution, we make use of the already introduced probability measure
$\mathbb{P}(M)=\frac{1}{n\nu_{n}}\int_{M\cap S^{n-1}}\!\,\mathrm{d}\eta$ (2)
on the Borel subsets $M$ of the $\mathbb{R}^{n}$. The direct calculation of the expected values of the given and other homogeneous functions as surface integrals is complicated. We therefore reduce the calculation of such quantities to the calculation of simpler volume integrals. Our main tool is the radial-angular decomposition (13).

### v5

@@ 1 Introduction
The numerical solution of partial differential equations in high space dimensions is a difficult and challenging task. Methods such as finite elements, which work perfectly in two or three dimensions, are not suitable for solving such problems because the effort grows exponentially with the dimension. Random walk based techniques only provide solution values at selected points. Sparse grid methods are best suited for problems in still moderate dimensions. Tensor-based methods Bachmayr , Hackbusch_1 , Khoromskij stand out in this area. They are not subject to such limitations and perform surprisingly well in a large number of cases. Tensor-based methods exploit the structure of the solution rather than its regularity. Consider the equation
@@
$U_{k+1}=\,U_{k}\,-\,(-\Delta+\mu)^{-1}(\mathcal{L}U_{k}+\mu U_{k}-F)$ (8)
starting from $U_{0}=0$. The calculation of the iterates requires the solution of equations of the form (1), that is, the calculation of integrals of the form (2), now over the higher dimensional $\mathbb{R}^{n}$. The symbol $\|T^{t}\omega\|^{2}$ of the operator $\mathcal{L}$ is a homogeneous second-order polynomial in $\omega$. For separable right-hand sides $F$ as above, the calculation of these integrals thus reduces to the calculation of products of lower, in the extreme case one-dimensional integrals.
The reason for the usually astonishingly fast convergence of this iteration is the directional behavior of the term $\|T^{t}\omega\|^{2}$. The higher the dimension $m$, the lower the probability that the values $\|T^{t}\eta\|^{2}$ for $\eta$ on the unit sphere $S^{n-1}$ of the $\mathbb{R}^{n}$ deviate much from their mean, a typical concentration of measure effect. To capture this phenomenon quantitatively, we introduce the probability measure
@@
on the Borel subsets $M$ of the $\mathbb{R}^{n}$, where $\nu_{n}$ is the volume of the unit ball and $n\nu_{n}$ thus is the area of the unit sphere. If $M$ is a subset of the unit sphere, $\mathbb{P}(M)$ is equal to the ratio of the area of M to the area of the unit sphere. If $M$ is a sector, that is, if $M$ contains with a vector $\omega$ also its scalar multiples, $\mathbb{P}(M)$ measures the opening angle of $M$. The quantity on which all our analysis is based is the angular distribution
$F(\delta)={\mathbb{P}\big{(}\big{\{}\omega\in\mathbb{R}^{n}\,\big{|}\,\|T^{t}\omega\|^{2}\leq\delta\|\omega\|^{2}\big{\}}\big{)}}$ (10)
of the values $\|T^{t}\omega\|^{2}$. We assume that its expected value, the mean value of the expression $\|T^{t}\eta\|^{2}$ over the unit sphere, is one. This is only a matter of the scaling of the variables in the higher dimensional space and does not represent a restriction. Apart from extreme cases, the distribution (10) approaches a normal distribution with increasing dimensions. The concentration of measure effect is reflected in the fact that the variance of the distribution decreases and tends, under rather general circumstances, to zero as the dimensions increase.
@@
with respect to a broad range of Fourier-based norms. This is admittedly an idealized situation and the actual convergence behavior is more complicated. Nevertheless, this example accurately describes what to expect. Provided that for points $\eta$ on the unit sphere the values $\|T^{t}\eta\|^{2}$ are approximately normally distributed with expected value $E=1$ and small variance $V$, the measure (9) of the sector $S$ is almost one as soon as $\rho$ exceeds the standard deviation $\sigma=\sqrt{V}$ by more than a moderate factor. The sector $S$ then fills almost the entire frequency space. The higher the dimensions and the smaller the variance, the faster the iterates approach the solution.
@@
In a sense, the following section forms the core of the present work. It is devoted to the study of the angular distribution (10) of the values $\|T^{t}\omega\|^{2}$. Our first result, Theorem 3.1, is a semi-explicit representation of the density of this distribution in form of an integral over the unit sphere in $\mathbb{R}^{m}$. It gives detailed information about the behavior of this distribution for small $\delta$. In a special case, when all singular values of $T$ are equal, it becomes a rescaled beta distribution. Moreover, we show how the expected value $E$ and the variance $V$ of the distribution can be expressed in terms of the singular values of the matrix $T$. Using this representation, we show that for random matrices $T$ with assigned expected value $E=1$ the expected value of the variances $V$ not only tends to zero as the dimension $m$ goes to infinity, but also that these variances cluster the more around their expected value the larger $m$ is. We also study a class of matrices $T$ that are associated with interaction graphs. These matrices correspond to the case that some or all coordinate differences are introduced as additional variables and formed the motivation for the present work. The expected values that are assigned to these matrices take the value one and the variances $V$ can be expressed directly in terms of the vertex degrees. Finally, we show how to efficiently sample the values $\|T^{t}\eta\|^{2}$ for a large number of points $\eta$ that are uniformly distributed on the unit sphere. Calculations of this kind support our claim that these values are approximately normally distributed in high dimensions.
In the final section, we return to the higher-dimensional counterpart (7) of the original equation (1) and examine the convergence behavior of the polynomially accelerated version of the iteration (8) for its solution. Special attention is given to the limit case $\mu=0$ of the Laplace equation. The section ends with a brief review of an approximation of the form (4) by sums of Gauss functions.
@@ 2 Solutions as traces of higher-dimensional functions
$\|U\|_{0}=\int|\widehat{U}(\omega)|\,\mathrm{d}\omega.$ (2)
@@
$(\mathrm{D}^{\alpha}u)(x)=\left(\frac{1}{\sqrt{2\pi}}\right)^{n}\!\int(\mathrm{i}\,T^{t}\omega)^{\alpha}\widehat{U}(\omega)\,\mathrm{e}^{\,\mathrm{i}\,\omega\cdot Tx}\,\mathrm{d}\omega$ (5)
that, like $u$, is itself uniformly continuous and vanishes at infinity.
@@
$\|U\|_{2}=\int\|T^{t}\omega\|^{2}\,|\widehat{U}(\omega)|\,\mathrm{d}\omega.$ (6)
@@
From now on, the equation (9) will replace the original equation (11). Our aim is to compute its solution (10) iteratively by polynomial accelerated versions of the basic iteration (8). The convergence properties of this iteration depend decisively on the directional behavior of the values $\|T^{t}\omega\|^{2}$, which will be studied in the next section, before we return to the equation and its iterative solution.
@@
The norms split into two groups and are labeled by a smoothness parameter. Like the differential operator (7) itself, they depend on the matrix $T$ under consideration. In a first step, we assign the function $\phi:S^{n-1}\to\mathbb{R}$ with the values
$\phi(\eta)=n\nu_{n}\int_{0}^{\infty}\|T^{t}\eta\|^{s}\,|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r$ (15)
to a given integrable function $U:\mathbb{R}^{n}\to\mathbb{R}$ and a given real smoothness parameter $s$. The two norms of $U$ are then defined via
$\|U\|_{s}=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\phi(\eta)\,\mathrm{d}\eta,\quad|\!|\!|U|\!|\!|_{s}^{2}=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\phi(\eta)^{2}\,\mathrm{d}\eta.$ (16)
If the norm $|\!|\!|U|\!|\!|_{s}$ of $U$ is finite, then the norm $\|U\|_{s}$ is also finite. In this case, the estimate $\|U\|_{s}\leq|\!|\!|U|\!|\!|_{s}$ holds. This follows from the Cauchy-Schwarz inequality. In cartesian coordinates, the first of the two norms takes the simpler form
$\|U\|_{s}=\int_{\mathbb{R}^{n}}\|T^{t}\omega\|^{s}\,|\widehat{U}(\omega)|\,\mathrm{d}\omega.$ (17)
The traces $u(x)=U(Tx)$ of functions $U$ with finite norms $\|U\|_{0}$ and $\|U\|_{s}$ for a real number $s\geq 1$ possess continuous partial derivatives of order $k$, $k\leq s$, which can be bounded in terms of the then also finite norms $\|U\|_{k}$ and vanish at infinity.
A direct consequence of the definitions are the following regularity theorems.
Theorem 2.2
For right-hand sides $F$ with finite norms $\|F\|_{s}$ and $|\!|\!|F|\!|\!|_{s}$, respectively, the solution (10) of the equation (9) satisfies the estimates
$\|U\|_{s+2}\leq\|F\|_{s},\quad|\!|\!|U|\!|\!|_{s+2}\leq|\!|\!|F|\!|\!|_{s}$ (18)
for all smoothness parameters $s$, independent of $\mu\geq 0$. If $\mu=0$, equality holds.
This simply results from the Fourier representation of the solution and the estimate
$\frac{1}{r^{2}\|T^{t}\eta\|^{2}+\mu}\leq\frac{1}{\|T^{t}\eta\|^{2}}\,\frac{1}{r^{2}}.$ (19)
If $\mu>0$, the left-hand side can be estimated by $1/\mu$ instead. This leads to
Theorem 2.3
For right-hand sides $F$ with finite norms $\|F\|_{s}$ and $|\!|\!|F|\!|\!|_{s}$, respectively, the solution (10) of the equation (9) satisfies the estimates
$\|U\|_{s}\leq\frac{1}{\mu}\,\|F\|_{s},\quad|\!|\!|U|\!|\!|_{s}\leq\frac{1}{\mu}\,|\!|\!|F|\!|\!|_{s}$ (20)
for all smoothness parameters $s$, provided $\mu$ is strictly positive.
3 The angular distribution
The purpose of this section is a very detailed study of the angular distribution of the values $\|T^{t}\omega\|^{2}$, based on our previously introduced probability measure
$\mathbb{P}(M)=\frac{1}{n\nu_{n}}\int_{M\cap S^{n-1}}\!\,\mathrm{d}\eta$ (1)
on the Borel subsets $M$ of the $\mathbb{R}^{n}$ that for a sector measures its opening angle. In terms of this probability measure, the angular distribution of these values is
$F(\delta)={\mathbb{P}\big{(}\big{\{}\omega\in\mathbb{R}^{n}\,\big{|}\,\|T^{t}\omega\|^{2}\leq\delta\|\omega\|^{2}\big{\}}\big{)}}$ (2)
or, after restriction to the unit sphere $\|\eta\|=1$ itself,
$F(\delta)={\mathbb{P}\big{(}\big{\{}\eta\in S^{n-1}\,\big{|}\,\|T^{t}\eta\|^{2}\leq\delta\big{\}}\big{)}}.$ (3)
The direct calculation of such and similar quantities is difficult. We therefore reduce the calculation of corresponding integrals over the unit sphere to the calculation of simpler volume integrals. Our main tool is the radial-angular decomposition (13).
@@ 3 The angular distribution
If the function $\chi:\mathbb{R}^{n}\to\mathbb{R}$ is positively homogeneous of nonnegative degree $\ell$, the integral of $\chi$ over the unit sphere is equal to the volume integral
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\chi(\eta)\,\mathrm{d}\eta=C(\ell)\int_{\mathbb{R}^{n}}\chi(\omega)W(\omega)\,\mathrm{d}\omega,$ (4)
where the rotationally symmetric function $W=W_{n}$ is the normed Gauss function
$W(\omega)=\left(\frac{1}{\sqrt{\pi}}\right)^{n}\exp\big{(}-\|\omega\|^{2}\big{)},$ (5)
which splits into a product of one-dimensional functions, and the prefactor is
$C(\ell)=\frac{\Gamma(n/2)}{\Gamma((n+\ell)/2)}.$ (6)
@@
This follows immediately from the radial angular decomposition of volume integrals into an inner radial and an outer angular part, the identities
$\nu_{n}=\frac{2}{n}\,\frac{\pi^{n/2}}{\Gamma(n/2)},\quad\int_{0}^{\infty}t^{j}\mathrm{e}^{-t^{2}}\,\mathrm{d}t=\frac{1}{2}\,\Gamma\left(\frac{j+1}{2}\right),$
and the homogeneity $\chi(r\eta)=r^{\ell}\chi(\eta)$ of the function under consideration. ∎
Let $H$ be the Heaviside function, with the values $H(t)=0$ for $t<0$ and $H(t)=1$ for $t\geq 0$, and let $\chi(\omega,\delta)=H(\delta\|\omega\|^{2}-\|T^{t}\omega\|^{2})$. As $\chi(\omega,\delta)$ is homogeneous of degree zero as a function of $\omega$, the distribution (2) possesses the representation
$F(\delta)=\int_{\mathbb{R}^{n}}\chi(\omega,\delta)W(\omega)\,\mathrm{d}\omega$ (7)
by Lemma 2. It depends only on the singular values of the matrix $T$. This can be shown by means of the singular value decomposition of $T$ and the transformation theorem for volume integrals. Since the expression $\chi(\omega,\delta)$ is right-continuous as a function of $\delta$, the distribution is right-continuous by the dominated convergence theorem. For dimensions $n>m+2$, it has a representation
$F(\delta)=\int_{-\infty}^{\delta}f(t)\,\mathrm{d}t$ (8)
with a density $f$ that attains values $f(t)>0$ on the interval $0<t<\|T^{t}\|^{2}$ and $f(t)=0$ outside of it and shows a characteristic, dimension-dependent behavior near $t=0$.
Theorem 3.1
If $n-m>2$, the distribution (2) possesses the density function $f$ that vanishes for $t\leq 0$ and is, for arguments $t>0$, given by
$f(t)=\frac{1}{B(\alpha,\beta)}\frac{1}{m\nu_{m}}\int_{S^{m-1}}\!\frac{t^{\alpha-1}R(1-\,t\,\|\Sigma_{0}^{-1}\eta\|^{2})^{\beta-1}}{\det\Sigma_{0}}\,\,\mathrm{d}\eta,$ (9)
where $\Sigma_{0}=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{m})$ is the $(m\times m)$-diagonal matrix with the singular values of the matrix $T$ on its diagonal, the coefficients $\alpha$ and $\beta$ are
$\alpha=\frac{m}{2},\quad\beta=\frac{n-m}{2},$ (10)
the function $R$ takes the values $R(t)=\max(0,t)$, and $B$ is Euler’s beta function.
Proof
[…]
$a(x,\delta)^{2}=R\bigg{(}\frac{\|\Sigma_{0}x\|^{2}-\delta\|x\|^{2}}{\delta}\bigg{)}.$
Since $\|\Sigma_{0}x\|^{2}\leq\delta(\|x\|^{2}+\|y\|^{2})$ holds if and only if $\|y\|-a(x,\delta)\geq 0$, the distribution can be written as a double integral as follows
$F(\delta)=\int_{\mathbb{R}^{m}}\bigg{(}\int_{\mathbb{R}^{n-m}}H\big{(}\|y\|-a(x,\delta)\big{)}W_{n-m}(y)\,\mathrm{d}y\bigg{)}W_{m}(x)\,\mathrm{d}x.$
This results from its representation (7) and Fubini’s theorem. By (14), the inner integral reduces to the one-dimensional integral
$\frac{2}{\Gamma((n-m)/2)}\int_{0}^{\infty}H\big{(}r-a(x,\delta)\big{)}\mathrm{e}^{-r^{2}}r^{(n-m)-1}\,\mathrm{d}r.$
Introducing the function
$\phi(s)=\frac{2}{\Gamma((n-m)/2)}\int_{s}^{\infty}r^{(n-m)-1}\mathrm{e}^{-r^{2}}\,\mathrm{d}r,$
the distribution therefore takes the form
$F(\delta)=\int_{\mathbb{R}^{m}}\phi(a(x,\delta))W_{m}(x)\,\mathrm{d}x.$
Let $x\neq 0$. Because $a(x,t)=(\|\Sigma_{0}x\|^{2}/t-\|x\|^{2})^{1/2}>0$ for $0<t<\|\Sigma_{0}x\|^{2}/\|x\|^{2}$, the function $t\to\phi(a(x,t))$ is differentiable on this interval. There it has the derivative
$h(x,t)=\frac{a(x,t)^{(n-m)-2}\exp(-\,a(x,t)^{2})\|\Sigma_{0}x\|^{2}}{\Gamma((n-m)/2)\,t^{2}}.$
Because $n-m>2$, this derivative tends to zero as $t$ goes to $\|\Sigma_{0}x\|^{2}/\|x\|^{2}$ from the left. Because $a(x,t)=0$ for $\|\Sigma_{0}x\|^{2}/\|x\|^{2}\leq t$, $\phi(a(x,t))$ is therefore continuously differentiable as a function of $t$ on the entire positive real axis $t>0$ and its derivative has the above representation. The same applies in the case $x=0$. For $0<\delta_{0}<\delta$
$F(\delta)-F(\delta_{0})=\int_{\mathbb{R}^{m}}\bigg{(}\int_{\delta_{0}}^{\delta}h(x,t)\,\mathrm{d}t\bigg{)}W_{m}(x)\,\mathrm{d}x$
follows, or, after interchanging the order of integration,
$F(\delta)-F(\delta_{0})=\int_{\delta_{0}}^{\delta}f(t)\,\mathrm{d}t,$
where the density for arguments $t>0$ is given by
$f(t)=\int_{\mathbb{R}^{m}}h(x,t)W_{m}(x)\,\mathrm{d}x.$
Because the exponential terms partially cancel each other out, it reads explicitly
$f(t)=\int_{\mathbb{R}^{m}}\frac{R(\|Ax\|^{2}-\|x\|^{2})^{(n-m)/2-1}\|Ax\|^{2}}{\Gamma((n-m)/2)\,t}\;W_{m}(Ax)\,\mathrm{d}x$
in terms of the matrix $A=\Sigma_{0}/\sqrt{t}$, or, after a change of variables,
$f(t)=\int_{\mathbb{R}^{m}}\frac{R(\|x\|^{2}-\|A^{-1}x\|^{2})^{(n-m)/2-1}\|x\|^{2}}{\Gamma((n-m)/2)\,t\,|\det A|}\;W_{m}(x)\,\mathrm{d}x.$
Using Lemma 2, $|\det A|=t^{-m/2}\det\Sigma_{0}$, and the representation
$B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$
of the beta function in terms of the gamma function, this volume integral can finally be converted into the surface integral (9).
Since the function (9) is locally integrable and the distribution function (7) is right-continuous, one can let $\delta_{0}$ tend to zero in the relation above. This leads to
$F(\delta)-F(0)=\int_{0}^{\delta}f(t)\,\mathrm{d}t$
for $\delta>0$. Since the kernel of the $(m\times n)$-matrix $T^{t}$ is as an $(n-m)$-dimensional subspace a set of volume measure zero, $F(0)=0$ follows from the representation (7) of the distribution. This concludes the proof. ∎
A function $\eta\to h(\|T^{t}\eta\|^{2})$ is integrable over the unit sphere $S^{n-1}$ if and only if the function $t\to h(t)f(t)$ is integrable over the real axis. If this is the case,
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}h\big{(}\|T^{t}\eta\|^{2}\big{)}\,\mathrm{d}\eta=\int_{-\infty}^{\infty}h(t)f(t)\,\mathrm{d}t$ (11)
holds and the high-dimensional integral over the unit sphere reduces to an integral in one space dimension. As follows from (9), the quantities
$\frac{1}{\|T^{t}\eta\|^{s}},\quad s>0,$ (12)
are therefore integrable over the unit sphere if and only if $s$ is less than $m$.
Figure 1: The probability distributions assigned to the densities (13) for $m=2^{k}$, $k=1,\ldots,16$, and $n=2m$
If the matrix $\Sigma_{0}=\sigma I_{m}$ is a scalar multiple the identity matrix, the integral (9) can be evaluated explicitly. In this case, the density is
$f(t)=\frac{1}{\sigma^{2}}\,g\Big{(}\frac{t}{\sigma^{2}}\Big{)},\quad g(t)=\frac{1}{B(\alpha,\beta)}\;t^{\alpha-1}(1-t)^{\beta-1},$ (13)
on the interval $0<t<\sigma^{2}$ and vanishes outside this interval. That is, the distribution is a rescaled variant of a beta distribution. For high dimensions, it therefore behaves almost like a normal distribution. Figure 1 shows the assigned distributions for the cases $\sigma^{2}=n/m$, $m=2^{k}$, $k=1,\ldots,16$, and $n=2m$. The higher the dimensions are, the more the distributions approach the step function that jumps at $\delta_{0}=1$ from zero to one and the more the values $\|T^{t}\eta\|^{2}$, $\|\eta\|=1$, cluster around one.
Something similar can be observed in the general case. This is reflected in the variances of the distributions. They tend with very high probability to zero as the dimensions increase. The expected value and the variance of the distributions are
$E=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\|T^{t}\eta\|^{2}\,\mathrm{d}\eta,\quad V=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\big{(}\|T^{t}\eta\|^{2}-E\big{)}^{2}\,\mathrm{d}\eta$ (14)
and play a fundamental role in the further considerations.